# aDipartimento di Fisica, Università di Milano - Bicocca I-20126 Milano, Italy

### v1

@@
A measure on the space of CFTs and pure 3D gravity
Alexandre Belina,b,c, Alexander Maloneyd,e,f, Florian Seefeldc,g,h
@@ 1 Introduction
The construction of the ensemble can be approached from multiple view-points. One can either define a set of moments for the CFT data (scaling dimensions, spin, and OPE coefficients of the CFT operators), and work in a moment expansion. This was the approach taken in [16], and the leading order statistics for the OPE coefficients is Gaussian. This reproduces many on-shell Euclidean wormholes of the bulk theory. The issue is that the moments of OPE coefficients are highly non-Gaussian [20, 21], and it quickly becomes hard to track the moments of the CFT data. Alternatively, one can follow the approach of [22] and define directly a matrix-tensor model whose potential is obtained from the crossing equation, which guarantees that all the moments are correctly encoded. The tensor model then generates, through its Feynman diagrams, the sum over bulk topologies [23], with a contribution for each topology that agrees with the bulk partition function on a given topology (even at finite central charge), formulated most elegantly as a Virasoro TQFT [24, 17] or conformal Turaev-Viro theory [25, 26]. The tensor model also encodes non-trivial statistics for the density of states, which in the bulk come from an off-shell wormhole with torus boundaries [27].
@@
where $C(c_{0},\epsilon_{c})$ represents the space of CFTs in some window of size $\epsilon_{c}$ around a mean central charge $c_{0}$ and $\cdot$ is any observable that can be defined over the ensemble (for example a torus partition function). This requires defining a normalizable measure on the space of CFTs, which will be one of our results. We will then investigate the extent to which such an ensemble can be dual to pure gravity in AdS3.
@@ 2 A measure on the space of CFTs
Within a given family of CFTs living on the same conformal manifold, there is a natural measure induced by the Zamolodchikov metric. However, there is no standard measure on the space of isolated CFTs, or no standard measure to compare isolated CFTs to families living on conformal manifolds. [^7] See [37] for a discussion of this question along with various propositions for measures on the space of all (even disconnected) CFTs. Since we are looking for a maximum ignorance ensemble, the most natural measure on the space of CFTs is to weigh evenly all isolated CFTs, and weigh the continuous families by the volume of their conformal manifolds. This gives the following measure over the space of CFTs in $C(c_{0},\epsilon_{c})$
@@ 2.1 CFTs with $c\leq 1$
The divergence is slower than that coming from the minimal models. This means that the minimal models dominate the space of CFTs in the ensemble below c=1c=1.999It is interesting to try and interpret the limit of the minimal models as a genuine CFT at $c=1$, which would have a continuous spectrum. See [39] for a discussion of this, or [40, 41] for the supersymmetric case. Since the measure is normalizable at fixed $\Delta_{\min}$, an average over theories can be performed. This was studied for example in [42], in particular the spectral form factor of this ensemble. We now discuss the space of CFTs at higher values of $c$.
@@ 4.1 Subgroup number and conjugacy classes
It is useful to mention some preliminaries about what exactly we mean by counting subgroups here, and some known results from the mathematical litterature. The number of subgroups of $S_{N}$, denoted $s(S_{N})$, traditionally counts all possible ways to construct a subgroup of $S_{N}$ using its elements. Therefore, if the group $G_{N}$ has several different embeddings in $S_{N}$, $s(S_{N})$ will count each of these as different subgroups. To date, the best known bounds on $s(S_{N})$ are found in [65, 66, Corollary 3.3 and Theorem 4.2 respectively], [^12] We are using the little $o$ notation, where $f(x)=o(g(x))$ is defined as $\left|f(x)\right|\leq\epsilon g(x)$ for any $\epsilon>0$ as $x\rightarrow\infty$. In other words, $f(x)=o(g(x))$ effectively means that $f$ grows much slower than $g$.
@@
However, this is the total number of subgroups, which discriminates among different embeddings of a same group $G_{N}$. Some of these might have similar or different numbers of orbits when acting on $k$-tuples. If we look at the groups $\{1,(12)\},\{1,(13)\}$ and $\{1,(12)(34)...(N-1\,N)\}$, these are all different embeddings of the cyclic group $\mathbb{Z}_{2}$. When acting on the 1-tuples $\{i\},i=1,...,N$, we can nevertheless see that both $\{1,(12)\}$ and $\{1,(13)\}$ have N1N-1 orbits, while $\{1,(12)(34)...(N-1\,N)\}$ has $N/2$ orbits. Therefore, when studying the set of subgroups of $S_{N}$ for orbifold purposes, what we are actually interested in is the number of subgroups with similar group action. Such subgroups (in their permutation group embedding) are called permutation isomorphic, which has been shown to be equivalent to the subgroup embeddings being conjugate [67], and we are thus interested in the number of subgroups up to conjugacy, which we will denote $c(S_{N})$. As mentioned in [66, Section 6], and by noting that there can be at most $N!$ subgroups in a same conjugacy class, we have similar asymptotics as 4.1 for the number of subgroup conjugacy classes, as we have
@@ 4.2 The transitivity and oligomorphy restrictions
Finally, we note that there is also a discussion of “typical” subgroups of $S_{N}$ which can be found in the subgroup theory literature, even if only at the state of a conjecture. From [66, Sections 4 and 6], the “typical” subgroups of $S_{N}$ are conjectured to be elementary abelian 2-groups, which have the form $(\mathbb{Z}_{2})^{q}$ for $q\in\{1,2,...,2/n\}$.There is a large number of these subgroups, and in fact this large number of elementary abelian 2-groups gives rise to the lower bound of 4.1. If these groups really saturate the space of subgroups of $S_{N}$, then a typical subgroup has $\sim N/2$ orbits on 1-tuples. This would immediately imply that most subgroups are not oligomorphic, and in fact the density of oligomorphic subgroups of $S_{N}$ would therefore also decay as $e^{-N^{2}}$ as $N\rightarrow\infty$, as expected.
@@ Acknowledgements
We are happy to thank Connor Behan, Nathan Benjamin, Suzanne Bintanja, Jose Calderon-Infante, Scott Collier, Jan de Boer, Lorentz Eberhardt, Victor Gorbenko, Thomas Hartman, Christoph Keller, Diego Liska, Miguel Montero, Kyriakos Papadodimas, Eric Perlmutter, Boris Post, Julian Sonner and Irene Valenzuela for stimulating discussions. The work of FS was supported by the Laboratory for Theoretical Fundamental Physics (LTFP) at the Ecole Polytechnique Federale de Lausanne, as well as the research grant 00025445 from Villum Fonden and by the Sapere Aude: DFF-Starting Grant 4251-00029B. This work was made possible by Institut Pascal at Université Paris-Saclay with the support of the program “Investissements d’avenir” ANR-11-IDEX-0003-01. Research of AM is supported in part by the Simons Foundation Grant No. 12574 and the Natural Sciences and Engineering Research Council of Canada (NSERC), funding reference number SAPIN/00047-2020
@@ Appendix B Lower-bounding the density of known oligomorphic and transitive families
When looking at groups in $\mathbb{O}_{N}$, we see that for a given $N$, there are as many possible groups as there are unordered partitions of $N$ into $N=n_{1}+n_{2}+...+n_{k}$, $k$ finite, given by $p(k,N)$, with the asymptotics given by [76, Theorem 4.2.1]:

### v2

@@
We define a normalizable measure on the space of two-dimensional conformal field theories, which we interpret as a maximum ignorance ensemble. We test whether pure quantum gravity in AdS3 is dual to the average over this ensemble. We find a negative answer, which implies that CFTs with a primary gap of order the central charge are highly atypical in our ensemble. We provide evidence that more generally, holographic CFTs are atypical in the space of all CFTs by finding similar results for permutation orbifolds: subgroups of $S_{N}$ with a good large $N$ limit are very sparse in the space of all subgroups. Along the way, we derive several new results on the space of CFTs. Notably we derive an upper bound on the spacing in central charge between CFTs, which is doubly exponentially small in the large central charge limit.
@@ 1 Introduction
The construction of the ensemble can be approached from multiple view-points. One can either define a set of moments for the CFT data (scaling dimensions, spin, and OPE coefficients of the CFT operators), and work in a moment expansion. This was the approach taken in [16], and the leading order statistics for the OPE coefficients is Gaussian. This reproduces many on-shell Euclidean wormholes of the bulk theory. The issue is that the moments of OPE coefficients are highly non-Gaussian [20, 21], and it quickly becomes hard to track the moments of the CFT data. Alternatively, one can follow the approach of [22] and define directly a matrix-tensor model whose potential is obtained from the crossing equation, which guarantees that all the moments are correctly encoded. The tensor model then generates, through its Feynman diagrams, the sum over bulk topologies [23], with a contribution for each topology that agrees with the bulk partition function on a given topology (even at finite central charge), formulated most elegantly as a Virasoro TQFT [24, 17] or conformal Turaev-Viro theory [25, 26]. The tensor model also encodes non-trivial statistics for the density of states, which in the bulk come from an off-shell wormhole with torus boundaries [27] (see also [28]).
@@
where $C(c_{0},\epsilon_{c})$ represents the space of CFTs in some window of size $\epsilon_{c}$ around a mean central charge $c_{0}$ and $\cdot$ is any observable that can be defined over the ensemble (for example a torus partition function). This requires defining a normalizable measure on the space of CFTs, which will be one of our results. We will then investigate the extent to which such an ensemble can be dual to pure gravity in AdS3.222Note that for rational CFTs, a sum over topologies in the dual TQFT can produce a measure over the space of CFTs with that central charge [29]. It would be interesting to understand the connection with this work.
@@ 2 A measure on the space of CFTs
Within a given family of CFTs living on the same conformal manifold, there is a natural measure induced by the Zamolodchikov metric. However, there is no standard measure on the space of isolated CFTs, or no standard measure to compare isolated CFTs to families living on conformal manifolds. [^8] See [39, 40] for a discussion of this question along with various propositions for measures on the space of all (even disconnected) CFTs. Since we are looking for a maximum ignorance ensemble, the most natural measure on the space of CFTs is to weigh evenly all isolated CFTs [^9] Another possibility is to consider as measure for individual CFTs the size of their discrete automorphism group. Picking this measure instead will not significantly affect the conclusions reached in this paper., and weigh the continuous families by the volume of their conformal manifolds. This gives the following measure over the space of CFTs in $C(c_{0},\epsilon_{c})$
@@ 2.1 CFTs with $c\leq 1$
The divergence is slower than that coming from the minimal models. This means that the minimal models dominate the space of CFTs in the ensemble below c=1c=1.111111It is interesting to try and interpret the limit of the minimal models as a genuine CFT at $c=1$, which would have a continuous spectrum. See [42] for a discussion of this, or [43, 44] for the supersymmetric case. Since the measure is normalizable at fixed $\Delta_{\min}$, an average over theories can be performed. This was studied for example in [45], in particular the spectral form factor of this ensemble. We now discuss the space of CFTs at higher values of $c$.
@@ 4.1 Subgroup number and conjugacy classes
It is useful to mention some preliminaries about what exactly we mean by counting subgroups here, and some known results from the mathematical litterature. The number of subgroups of $S_{N}$, denoted $s(S_{N})$, traditionally counts all possible ways to construct a subgroup of $S_{N}$ using its elements. Therefore, if the group $G_{N}$ has several different embeddings in $S_{N}$, $s(S_{N})$ will count each of these as different subgroups. To date, the best known bounds on $s(S_{N})$ are found in [68, 69, Corollary 3.3 and Theorem 4.2 respectively], [^14] We are using the little $o$ notation, where $f(x)=o(g(x))$ is defined as $\left|f(x)\right|\leq\epsilon g(x)$ for any $\epsilon>0$ as $x\rightarrow\infty$. In other words, $f(x)=o(g(x))$ effectively means that $f$ grows much slower than $g$.
@@
However, this is the total number of subgroups, which discriminates among different embeddings of a same group $G_{N}$. Some of these might have similar or different numbers of orbits when acting on $k$-tuples. If we look at the groups $\{1,(12)\},\{1,(13)\}$ and $\{1,(12)(34)...(N-1\,N)\}$, these are all different embeddings of the cyclic group $\mathbb{Z}_{2}$. When acting on the 1-tuples $\{i\},i=1,...,N$, we can nevertheless see that both $\{1,(12)\}$ and $\{1,(13)\}$ have N1N-1 orbits, while $\{1,(12)(34)...(N-1\,N)\}$ has $N/2$ orbits. Therefore, when studying the set of subgroups of $S_{N}$ for orbifold purposes, what we are actually interested in is the number of subgroups with similar group action. Such subgroups (in their permutation group embedding) are called permutation isomorphic, which has been shown to be equivalent to the subgroup embeddings being conjugate [70], and we are thus interested in the number of subgroups up to conjugacy, which we will denote $c(S_{N})$. As mentioned in [69, Section 6], and by noting that there can be at most $N!$ subgroups in a same conjugacy class, we have similar asymptotics as 4.1 for the number of subgroup conjugacy classes, as we have
@@ 4.2 The transitivity and oligomorphy restrictions
Finally, we note that there is also a discussion of “typical” subgroups of $S_{N}$ which can be found in the subgroup theory literature, even if only at the state of a conjecture. From [69, Sections 4 and 6], the “typical” subgroups of $S_{N}$ are conjectured to be elementary abelian 2-groups, which have the form $(\mathbb{Z}_{2})^{q}$ for $q\in\{1,2,...,2/n\}$.There is a large number of these subgroups, and in fact this large number of elementary abelian 2-groups gives rise to the lower bound of 4.1. If these groups really saturate the space of subgroups of $S_{N}$, then a typical subgroup has $\sim N/2$ orbits on 1-tuples. This would immediately imply that most subgroups are not oligomorphic, and in fact the density of oligomorphic subgroups of $S_{N}$ would therefore also decay as $e^{-N^{2}}$ as $N\rightarrow\infty$, as expected.
@@ Acknowledgements
We are happy to thank Connor Behan, Nathan Benjamin, Suzanne Bintanja, Jose Calderon-Infante, Scott Collier, Jan de Boer, Tolya Dymarsky, Lorentz Eberhardt, Victor Gorbenko, Thomas Hartman, Christoph Keller, Diego Liska, Miguel Montero, Kyriakos Papadodimas, Eric Perlmutter, Boris Post, Julian Sonner and Irene Valenzuela for stimulating discussions. The work of FS was supported by the Laboratory for Theoretical Fundamental Physics (LTFP) at the Ecole Polytechnique Federale de Lausanne, as well as the research grant 00025445 from Villum Fonden and by the Sapere Aude: DFF-Starting Grant 4251-00029B. This work was made possible by Institut Pascal at Université Paris-Saclay with the support of the program “Investissements d’avenir” ANR-11-IDEX-0003-01. Research of AM is supported in part by the Simons Foundation Grant No. 12574 and the Natural Sciences and Engineering Research Council of Canada (NSERC), funding reference number SAPIN/00047-2020
@@ Appendix B Lower-bounding the density of known oligomorphic and transitive families
When looking at groups in $\mathbb{O}_{N}$, we see that for a given $N$, there are as many possible groups as there are unordered partitions of $N$ into $N=n_{1}+n_{2}+...+n_{k}$, $k$ finite, given by $p(k,N)$, with the asymptotics given by [79, Theorem 4.2.1]: